Determine the quadratic equation with integral coefficients having the roots:
-3 and 1
: Determine the quadratic equation with integral coefficients having the roots:
-3 and 1

Answer by checkley77(3654)

X^2+2X-3=0
(X+3)(X-1)=0
X+3=0
X=-3 ANSWER.
X-1=0
X=1 ANSWER.